# Generating Plausible Distractors for Multiple-Choice Questions via Student Choice Prediction

### v1

@@ B.8 Error Analysis
B.9 GPT Prompt for Distractor Validity Check
The instruction prompt for checking validity of distractors is in Table 25. If the output is ‘invalid’ (as it is an incorrect option for the question), it is considered a distractor.
@@ Appendix C Base MCQ Dataset
Appendix D GPT Prompt for Augmenting Distractors in Base MCQ Dataset
The instruction prompt for augmenting distractors in base MCQ dataset is in Table 26. Through this prompt, the student choice dataset was constructed only when at least one newly generated distractor by GPT-4o was valid and did not overlap with the original.
@@ Appendix E Potential Issues
  Prompt for Making Pairwise Ranker Training Data
@@
Table 18: Intruction prompt for making pairwise ranker training data.
@@
  Prompt for Distractor Validity Check
@@
Table 25: Instruction prompt for distractor validity check.
  Prompt for Augmenting Distractors in Base MCQ Dataset
@@
Table 26: Instruction prompt for augmenting distractors in base MCQ dataset.

### v2

@@ B.7 Error Analysis
B.8 Prompt for Checking Distractor Validity
The instruction prompt for checking the validity of distractors is in Table 29. If the output is ‘invalid’ (as it is an incorrect option for the question), it is considered a distractor.
@@ Appendix C Base MCQ Dataset
Appendix D Prompt for Augmenting Distractors in the Base MCQ Dataset
The instruction prompt for augmenting distractors in the base MCQ dataset is in Table 30. Through this prompt, the student choice dataset was constructed only when at least one newly generated distractor by GPT-4o was valid and did not overlap with the original.
@@ Appendix E Potential Issues
  Prompt for Generating Pairwise Ranker Training Data
@@
Table 22: Instruction prompt for generating pairwise ranker training data.
@@
  Prompt for Checking Distractor Validity
@@
Table 29: Instruction prompt for checking the validity of distractors.
  Prompt for Augmenting Distractors in the Base MCQ Dataset
@@
Table 30: Instruction prompt for augmenting distractors in the base MCQ dataset.